I need help determining whether each pair of equations represents perpendicular lines.
y = –x + 7, y – x = 3
Thank you so much!

Answer by sudhanshu_kmr(1152)


y = -x +7
slope of first equation = -1
y - x = 3
=> y = x+3
slope of second equation = 1
product of slopes = 1 * -1 = -1
so, these are perpendicular lines.